# Hierarchical Multi-Agent MCTS for Safety-Critical Coordination in Mixed-Autonomy Roundabouts

### v3

@@ V-B Case 2: Dual-lane Roundabout (ROP = 50%)
Figure 10: PET distributions and violations under different CAV penetration rates.
@@
As presented in Table III, our MCTS-based method with $d_{\text{max}}=8$ demonstrates consistently strong performance in mixed traffic environments. It achieves a high arrival rate of $92.9\pm 3.2\%$ alongside a remarkably low collision rate of just $1.5\pm 0.5\%$. Although the average computation time ($30.1\pm 5.0$ s) is slightly higher than that of the baseline method ($25.1\pm 3.5$ s), the trade-off is justified by the significant gains in safety and efficiency. In contrast, traditional approaches such as Stackelberg and Nash suffer from considerably higher collision rates ($33.0\%$ and $28.5\%$, respectively) and much lower arrival rates, underscoring the limitations of these methods in the presence of unpredictable human-driven vehicles.
@@ VI Conclusion
This paper introduces a safety-critical decision-making framework for autonomous vehicles navigating unsignalized, dual-lane roundabouts by integrating Monte Carlo Tree Search (MCTS) with a hierarchical risk assessment strategy. The framework offers three major innovations: a multi-agent MCTS structure for scalable and efficient action space exploration, a hierarchical safety assessment mechanism for robust spatiotemporal risk evaluation, and an adaptive reward function that effectively balances safety and efficiency. Experimental results confirm the effectiveness of the proposed method under varying autonomous vehicle penetration rates. In fully autonomous settings ($100\%$ CAVs), the framework achieves reduced trajectory deviations and eliminates PET violations when compared to baseline approaches. In mixed traffic scenarios ($50\%$ CAVs + $50\%$ HDVs), the framework delivers even greater improvements by reliably handling the uncertainty introduced by human drivers, while maintaining low deviation and high safety margins. These results highlight the framework’s strong potential for real-world deployment in complex traffic environments. Future work will focus on improving computational scalability, including the development of pruning techniques and parallelized search strategies to mitigate the combinatorial growth of the action space. Furthermore, extending the framework to accommodate diverse roundabout geometries will enhance its applicability across broader real-world contexts.

### v4

@@ V-B Case 2: Dual-lane Roundabout (ROP = 50%)
Figure 10: PET violations under different CAV penetration rates.
@@
As shown in Table V, our MCTS-based method with $d_{\max}=8$ achieves the best overall performance in Case 2, yielding the lowest collision rate (2.0%) and the smallest trajectory deviation ($0.63\pm 0.59$,m), while maintaining a high average speed ($5.23\pm 0.94$,m/s). In comparison, Stackelberg and Nash exhibit substantially higher collision rates (33.0% and 29.0%), highlighting their limited robustness in mixed traffic with human-driven uncertainty. These results confirm that increasing $d_{\max}$ significantly improves safety and motion consistency in complex roundabout interactions.
@@
Figure 11: Ablation results. (a) Reward allocation strategies. (b) HDV uncertainty models.
TABLE VI: Sensitivity and ablation analysis
Setting Collision PET Viol. Traj. Dev. Min.
Rate (%) Rate (%) (m) Distance (m)
Min. Safe Distance $d_{\min}$ (Default $d_{\min}=2.0$ m)
$d_{\min}=1.0$ m 9.0 8.2 $\pm$ 0.9 0.56 $\pm$ 0.39 1.31 $\pm$ 0.74
$\bm{d_{\min}=2.0}$m 2.0 3.1 $\pm$ 0.7 0.73 $\pm$ 0.61 2.72 $\pm$ 0.93
$d_{\min}=3.0$ m 1.0 2.4 $\pm$ 0.5 1.42 $\pm$ 0.85 3.61 $\pm$ 1.12
Reward Allocation Strategy (Default Shapley)
Equal sharing 14.0 10.2 $\pm$ 2.1 1.63 $\pm$ 0.85 1.12 $\pm$ 0.71
Distance-weighted 7.0 5.3 $\pm$ 1.6 1.12 $\pm$ 0.78 2.36 $\pm$ 0.87
Shapley value 2.0 3.1 $\pm$ 0.7 0.73 $\pm$ 0.61 2.71 $\pm$ 0.93
HDV Uncertainty Model (Default Lane-Specific)
Deterministic IDM 0.0 1.9 $\pm$ 0.5 0.52 $\pm$ 0.43 2.54 $\pm$ 0.98
Lane-agnostic 11.0 8.7 $\pm$ 1.8 1.66 $\pm$ 0.82 1.25 $\pm$ 0.67
Lane-specific 2.0 3.1 $\pm$ 0.7 0.73 $\pm$ 0.61 2.71 $\pm$ 0.93
V-C Sensitivity and Ablation Analysis
To evaluate the robustness of key design choices and validate the contributions of individual components, we conduct sensitivity and ablation studies. The quantitative results are summarized in Table VI, with visual comparisons for the two core algorithmic contributions presented in Fig. 11.
V-C1 Sensitivity to Minimum Safe Distance
We vary the minimum safe distance threshold $d_{\min}\in\{1.0,2.0,3.0\}$ m. As shown in Table VI, a smaller threshold ($d_{\min}=1.0$ m) permits closer interactions, resulting in elevated collision rates ($9.0\%$) and reduced minimum distances ($1.31$ m). Conversely, a larger threshold ($d_{\min}=3.0$ m) yields slightly lower collision rates ($1.0\%$) but introduces excessive conservatism, increasing trajectory deviations to $1.42$ m. The default setting $d_{\min}=2.0$ m achieves an effective balance between safety and efficiency, maintaining a low collision rate ($2.0\%$) with moderate trajectory deviations ($0.73$ m).
V-C2 Reward Allocation Strategy
We compare three reward allocation strategies: equal sharing, distance-weighted sharing, and Shapley value allocation. As illustrated in Fig. 11(a), the Shapley value method significantly outperforms both alternatives. Equal sharing, which ignores individual contributions, leads to the highest collision rate ($14.0\%$) and PET violation rate ($10.2\%$), as agents lack incentive for cooperative behavior. Distance-weighted sharing improves upon equal sharing but still yields a $7.0\%$ collision rate. The Shapley value method reduces collision rates to $2.0\%$ and achieves the best minimum distance ($2.71$ m), demonstrating that fairly attributing each agent’s marginal contribution to the coalition reward is essential for effective multi-agent coordination.
V-C3 HDV Uncertainty Model
We evaluate three uncertainty modeling approaches for human-driven vehicles: deterministic IDM, lane-agnostic uncertainty, and lane-specific uncertainty. As shown in Fig. 11(b), the deterministic IDM achieves zero collisions in simulation, which is expected since the prediction model matches the simulator. However, this zero-margin approach offers no robustness against real-world behavioral deviations. The lane-agnostic model, which applies uniform uncertainty across lanes, introduces excessive conservatism that paradoxically increases collisions to $11.0\%$, as overly cautious planning causes hesitation and deadlock. Our lane-specific model strikes an effective balance, reducing collisions to $2.0\%$ while providing structured robustness margins calibrated to lane-dependent behavioral variability, as outer-lane vehicles exhibit inherently higher trajectory uncertainty than inner-lane vehicles.
@@ VI Conclusion
This paper introduces a safety-critical decision-making framework for autonomous vehicles navigating unsignalized, dual-lane roundabouts by integrating Monte Carlo Tree Search (MCTS) with a hierarchical risk assessment strategy. The framework offers three major innovations: a multi-agent MCTS structure for scalable and efficient action space exploration, a hierarchical safety assessment mechanism for robust spatiotemporal risk evaluation, and an adaptive reward function that effectively balances safety and efficiency. Experimental results confirm the effectiveness of the proposed method under varying autonomous vehicle penetration rates. In fully autonomous settings ($100\%$ CAVs), the framework achieves reduced trajectory deviations and eliminates PET violations when compared to baseline approaches. In mixed traffic scenarios ($50\%$ CAVs + $50\%$ HDVs), the framework delivers even greater improvements by reliably handling the uncertainty introduced by human drivers, while maintaining low deviation and high safety margins. Sensitivity and ablation analyses further demonstrate robustness to key parameter choices and confirm that each proposed component—Shapley value allocation and lane-specific uncertainty modeling—contributes meaningfully to overall performance. We note that the current framework assumes accurate state observations and plans independently per CAV without inter-vehicle communication. While these assumptions are idealized, the probabilistic C2H uncertainty model provides inherent robustness margins that can partially absorb sensing inaccuracies. The current results, validated in simulation, demonstrate the framework’s promise as an effective decision-support tool for complex roundabout coordination. Future work will focus on improving computational scalability through parallelized search strategies, extending the framework to diverse roundabout geometries, and integrating data-driven behavior prediction with validation under realistic sensing noise and communication delays.